Hey, I really need help in this question:
Joe can run twice as fast as Pete. They start at the same point and run in opposite direction for 40 minutes. The distance between them is now 16km. How fast does Joe run?
I need the working out steps
thanks

Answer by mananth(16946)

Joe can run twice as fast as Pete. They start at the same point and run in opposite direction for 40 minutes. The distance between them is now 16km. How fast does Joe run?
pete speed be x km/h
joe speed = 2x km/h
...
time = 40/60 = 2/3 hours
..
distance 16 km
..
since they are running away from each othet the speed will be x+2x= 3x km/h
..
distance /speed = time
16/3x = 2/3
3*16/3*2= x
x= 8 km/h
joe's speed = 2x= 16km/h
...
m.ananth@hotmail.ca